40m QIL Cryo_Lab CTN SUS_Lab CAML OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Fri Oct 9 17:41:04 2009, Jenne, Update, PEM, Accelerometers relocated 
    Reply  Sun Oct 11 04:12:44 2009, rana, Update, PEM, Accelerometers relocated Untitled.png
       Reply  Mon Oct 19 17:12:54 2009, jenne, Update, PEM, accelerometers and seismomters are all good. SeismometersGoodSpectra_19Oct2009.png
Message ID: 2119     Entry time: Mon Oct 19 17:12:54 2009     In reply to: 2079
Author: jenne 
Type: Update 
Category: PEM 
Subject: accelerometers and seismomters are all good. 


Some of these channels are not like the others.

 All of the PEM channels seem to be okay right now.  The accelerometers didn't turn out to have any differences in the traces when we put both XYZ triplets right next to each other, so we put them back where they belong.  Gur2 seismometer was showing a few problems, especially with Gur2_X, as Rana posted in elog 2079.  This was solved by tightening the cable screws which hold the Dsub end of the Guralp cable to the front panel of the Guralp box.  All is now well.

Attachment 1: SeismometersGoodSpectra_19Oct2009.png  124 kB  | Hide | Hide all
ELOG V3.1.3-